# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Unified Approach for Elevating Benchmark Quality

### v2

@@
Abstract.
Benchmarks are fundamental for ensuring unified evaluation and reproducibility. The rapid adoption of Artificial Intelligence in Software Engineering (AI4SE) has led to numerous benchmarks for tasks such as code generation and bug repair. However, this proliferation has introduced major challenges: (1) fragmented knowledge of benchmarks across tasks, (2) difficulty in selecting contextually relevant benchmarks, (3) lack of standardization in benchmark creation and enhancement, and finally (4) flaws that limit the utility of benchmarks. Addressing these challenges requires a dual approach: first, systematically mapping the current benchmark ecosystem to enable informed selection, and second, establishing unified guidelines for developing robust and adaptable benchmarks.
In this work, we first provide a comprehensive review of 247 studies from which we identify 273 benchmarks for evaluating AI4SE tasks since 2014. We categorize these benchmarks and analyze their limitations to highlight gaps in current benchmarking practices.
Drawing on the insights derived from the review, we built an extensible semantic search tool, BenchScout, to locate suitable benchmarks. We utilized an automated clustering technique that incorporates embedded contexts of a given benchmark’s associated studies.
We then applied dimensionality reduction and performed a parameter search to automatically optimize the configuration. To evaluate BenchScout’s usability, effectiveness, and intuitiveness, we conducted a user study with 22 participants, yielding average scores of 4.5, 4.0, and 4.1 out of 5, respectively.
Next, to improve benchmarking standards, we propose BenchFrame, a unified approach to elevate benchmarks’ quality. As a case study, we applied BenchFrame to the widely-used HumanEval benchmark to identify and rectify its key shortcomings. This resulted in HumanEvalNext which includes (1) corrected errors, (2) better language conversion, (3) increased test coverage, and (4) increased difficulty. We then assessed 10 state-of-the-art code language models on HumanEval, HumanEvalPlus, and HumanEvalNext. Compared to HumanEval and HumanEvalPlus, models’ performance on HumanEvalNext drops by an average of $31.22\%$ and $19.94\%$ in the pass@1 scores, respectively. This significant drop illustrates the values of continuous improvement for higher quality benchmarks. Moreover, we tackle issues related to the scalability of BenchFrame through a case study involving an agentic pipeline that delivers similar results. Similarly, we confront questions of generalizability by employing BenchFrame on the MBPP dataset, which equally impacts models’ performance comparably.
Lastly, we publicly release the material of our review, user study, and the enhanced benchmark.  [^1] https://github.com/AISE-TUDelft/AI4SE-benchmarks
Benchmark, Code Generation, AI4SE, Large Language Models, Search
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
1. Introduction
@@
Figure 1. ChatGPT repeating the same error from HumanEval (Screenshot captured in Dec 2023)
A notable example of benchmark limitations in code generation evaluation is HumanEval (Chen et al., 2021), a widely-used dataset that has been used for assessing large language models (LLMs) such as Codex (Chen et al., 2021), Gemini (Gemini Team, 2023), and GPT-4 (OpenAI, 2023). HumanEval was downloaded 82 thousand times in July 2025 on a single platform (Hugging Face)  [^2] https://huggingface.co/datasets/openai/openai_humaneval, which supports the strong and sustained interest in this benchmark. Despite its broad adoption, HumanEval contains numerous flaws and inconsistencies (van Dam et al., 2024). For instance, Task 47, which requires computing the median of a numerical list, incorrectly states the median of the list [-10,4,6,1000,10,20] is $15$. When queried on this task, ChatGPT-3.5 Turbo reproduced the same incorrect result (see Figure 1). This suggests potential data contamination and benchmark overfitting which can artificially inflate performance scores.
As HumanEval remains a widely-used benchmark in both AI and software engineering communities, several efforts have sought to expand its language support (Athiwaratkun et al., 2022; Cassano et al., 2022; Muennighoff et al., 2023; Zheng et al., 2023) or improve test coverage (Dong et al., 2023; Liu et al., 2023a). However, these extensions often build upon the original dataset without addressing its fundamental deficiencies, allowing inherent issues—such as flawed canonical solutions, vague problem definitions, incorrect tests, and insufficient coverage—to persist. Moreover, LLM-augmented improvements, such as automatic translation, often lack rigorous quality control as well. Lastly, as models have advanced, HumanEval and similar popular benchmarks have become increasingly saturated, with close to 100% scores for recent models.  [^3] https://evalplus.github.io/leaderboard.html This necessitates continuous elevation of the complexity of programs to better reflect the capabilities of contemporary models.
@@
While BenchScout aids navigation, our analysis confirms that many foundational benchmarks are flawed, which means a systematic solution is required to deal with them. Based on the identified gaps and limitations in current benchmarks, we introduce BenchFrame, a unified approach to enhance the quality of both existing and new benchmarks. To demonstrate its efficacy, we conduct a case study on HumanEval and present HumanEvalNext as an enhanced version.
When evaluating performance using HumanEval (original) and HumanEvalNext (our improved version based on the BenchFrame), we observe a substantial decline in pass@1 accuracy. Across ten state-of-the-art open-source code models, the average pass@1 score decreases by 31.2%, with a median drop of 26.0%. Performance remains significantly lower even on HumanEvalPlus (Liu et al., 2023a), an enhanced version of HumanEval, with an average decline of $19.94\%$ in pass@1 scores. These results highlight the importance of continuously refining benchmarks to better guide future research and provide more realistic assessments of model performance.
@@
  • We conducted a comprehensive review of 247 studies, identified 273 AI4SE benchmarks, and analyzed their limitations and gaps (Section 3),
  • We developed and released BenchScout, an extensible semantic search tool to facilitate locating appropriate AI4SE benchmarks. Our user study with 22 participants demonstrated its effectiveness (Section 4),
  • We propose a unified approach, BenchFrame, to improve the quality and reliability of benchmarks. A case study on HumanEval resulted in a refined, peer-reviewed benchmark, HumanEvalNext (Section 5). We assessed ten recent LLMs on this benchmark and verified significant drops in performance compared to the original HumanEval(Chen et al., 2021) and the enhanced version, HumanEvalPlus (Liu et al., 2023a),
@@ 2. Research Questions
These research questions directly map to the three-stage approach of our study. RQ1 fulfills the diagnostic part of our objective by mapping the landscape and identifying its core problems. The findings from RQ1 directly motivate RQ2, which provides a practical tool to help the community navigate the benchmark fragmentation we uncovered. Finally, RQ3 addresses the root cause of the quality issues found in RQ1 by proposing and validating a systematic framework for improvement, thereby fulfilling the final part of our research objective.
3. Existing Benchmarks, a Review
3.1. Search Criteria and Quality Assessment
@@
We employed a systematic method to search for, identify, and classify AI4SE benchmarks. This procedure involved three primary stages: conducting structured searches, verifying credibility, and developing a taxonomy. We searched on two platforms, namely Google Scholar and Semantic Scholar using the following keywords: “Benchmarks”, “Software Engineering”, “Large Language Models”, “Evaluation”, and “AI4SE”. We additionally search for benchmarks present in the PapersWithCode datasets collection due to the popularity and wide usage of the platform  [^6] This platform has since been discontinued and no longer accessible. The previously used url was https://paperswithcode.com/datasets.
@@
We selected these keywords based on a preliminary assessment of highly-cited benchmark research. Three authors worked together to revise the search criteria to ensure they were both relevant and complete. Our selection criteria targeted primary studies published in English from 2014 to 2025. After the initial paper collection process, duplicates were removed. Two authors reviewed the relevant of the identified pieces of literature, during which the originality (that is, the status as primary study), reproducibility, and accessibility of each study was accessed.
@@
More specifically, for originality, we accessed whether a study was a primary study that introduced a new benchmark or a significant extension of an existing one. This filter was set in place to distinguish between benchmark-proposing papers from studies that merely applied them with a certain level preprocessing. in terms of Reproducibility and Accessibility, we evaluated whether the benchmark and its associated resources were made publicly available. We looked for the presence of a public repository (e.g., GitHub, HuggingFace, etc.) which would allow the community to access and use the benchmark. After this phase, we performed forward and backward snowballing. A total of 247 papers were found through this process.
@@
After the papers were collected, the authors worked together to develop and consistently improve a taxonomy to efficiently categorize the benchmarks and extract metadata. Through continuous discussions, we identified initial essential categories while systematically gathering additional details for each study, including DOI and publication date. An iterative strategy was used to design the taxonomy and categorize SE tasks, starting with overarching categories such as reasoning, synthesis, and debugging. When the categories became too broad or lacked cohesion, the authors refined them into more detailed subcategories, establishing a multi-tiered hierarchy.
@@
In our evaluation, we not only considered benchmark datasets from academic sources, but also included those suggested by industry and subsequently adopted by scholars (for example, IBM CODAIT and Aider (CODAIT, 2021)). Our replication package also serves as a dynamic repository, allowing researchers to contribute additional benchmarks and related papers by submitting a pull request that includes the paper’s DOI.
@@
3.2. Results of the Review
Figure 2. Number of published benchmark papers in the last decade. (Jan 2014, Aug 2025)
@@ 3.2. Results of the Review
Additionally, upon analyzing the distribution of the tasks included in the review, we find that a great portion of benchmarks are related to Code Generation, more specifically, 31.8% of all the benchmarks included. Along with Project-Level Code Generation/Analysis and Code Understanding (with respectively 18.7% and 12.0%), these make up more than 60% of all benchmarks included. This, in itself, underscores a lack of attention to tasks such as vulnerability detection and code retrieval/search (cf. Figure 3).
The data reveal a notable change over time in the distribution of research categories throughout the observed period (cf. Figure 4). A key insight is the shift from code repair and maintenance to automated synthesis, marked by two opposing trends. Initially dominant, the Repair & Maintenance category gradually experiences a notable decline in its share. In contrast, starting around 2020, the Code Generation category gains prominence and reaches its peak in 2022, which coincides with the advancement of large-scale language models. In the recent years presented, there is an emerging trend of thematic diversification; as the focus on Code Generation recedes from its high, categories like Code Understanding and Retrieval & Search show a relative resurgence. This suggests a community-wide transition towards tackling additional challenges related to verification, integration, and the understanding of algorithmically generated outputs.
In the following, we summarize key insights from our review.
Code Generation (31.8%) Project-Level Analysis (18.7%) Code Understanding (12.0%) Math & Data (10.7%) Repair & Maintenance (10.7%) Other (6.7%) Vulnerability Detection (5.4%) Retrieval & Search (4.0%)
Figure 3. The distribution of the high-level categories of the benchmarks included in the study.
Figure 4. The progression of task distribution for the benchmarks included in the study;
Table 1. Overview of AI4SE benchmarks stemming from HumanEval (Chen et al., 2021).
@@
Original HumanEval (Chen et al., 2021) Python Avg. 7.7
Improved Language Support MultiPL-HumanEval (Cassano et al., 2022) 18 programming languages Avg. 7.7
HumanEval-Fix (Muennighoff et al., 2023) 6 programming languages Avg. 7.7
HumanEval-Explain (Muennighoff et al., 2023) 6 programming languages Avg. 7.7
HumanEval-Synthesize (Muennighoff et al., 2023) 6 programming languages Avg. 7.7
HumanEval-X (Zheng et al., 2023) 5 programming languages Avg. 7.7
Multi-HumanEval (Athiwaratkun et al., 2022) 12 programming languages Avg. 7.7
HumanEvalXL (Peng et al., 2024) 12PLs, 23NLs Avg. 8.33
mHumanEval (Raihan et al., 2024) Python, 204NLs Avg. 7.7
Improved Testing HumanEval+ (Liu et al., 2023a) Python Scaled ×80
HumanEval-MINI (Liu et al., 2023a) Python Scaled ×47
HE-Eval (Dong et al., 2023) Python Scaled ×14
@@
Extended EvoEval (Xia et al., 2024) Python Multiple categories,
@@
HumanEval Benchmark family: Currently, one of the most popular AI4SE benchmarks is HumanEval (Chen et al., 2021), used to evaluate the performance of many notable code-aware models (e.g., Codex (Chen et al., 2021), Gemini (Gemini Team, 2023), and GPT-4 (OpenAI, 2023)). This benchmark is used mainly for code synthesis, though there also exist some variations for code repair and code explanation (Muennighoff et al., 2023). Table 1 presents the family of HumanEval benchmarks. After an in-depth analysis of these benchmarks, we identified the following issues: (1) incorrect tests, (2) lack of proper test coverage, (3) incorrect canonical solutions, and (4) imprecise problem definitions. While there are versions that have improved the language support (Athiwaratkun et al., 2022; Cassano et al., 2022; Muennighoff et al., 2023; Zheng et al., 2023) and test coverage (Dong et al., 2023; Liu et al., 2023a), there is no version that contains all the improvements combined nor fixed the original issues. The issues for enhancing the original dataset can be generalized as follows:
@@
Table 2. Overview of AI4SE benchmarks derived from MBPP (Austin et al., 2021).
@@
Original MBPP (Austin et al., 2021) Python 974
Improved Language Support MultiPL-MBPP (Cassano et al., 2022) 18 programming languages 354-397 per language
MBXP (Athiwaratkun et al., 2022) 13 programming languages 848-974 per language
Improved Testing MBPP+ (Liu et al., 2023a) Python 427
MBPP-Eval (Dong et al., 2023) Python 974
MBPP Benchmark Family: Another AI4SE benchmark, highly similar in style and popularity compared to HumanEval, is MBPP (Austin et al., 2021): Mostly Basic Python Problems. It contains nearly a thousand crowdsourced problems, where almost half of it is sanitized and separately released. Furthermore, several enhancements have been published for MBPP (Table 2). Upon a more in-depth analysis of MBPP and its family of benchmarks, there are many signs suggesting deficient quality. One notable problem is the lack of proper testing, as MBPP originally only has three (rather trivial) tests per problem – which are all revealed in the prompt as well. With such a test suite in place, evaluation metrics become unstable and insignificant for proper comparison. The strength of the written tests and solutions themselves is not only troublesome in the original data but also the sanitized data features many flaws (even in corrected variants (Liu et al., 2023a)). From negligible observations such as poor syntax (e.g., too many spaces, Python method names starting with a capital – this is a common convention to only use for classes) to uncaught bugs and edge cases that break the implementation. While there are enhancements that improve the language support and extend the test cases, they are all built upon inadequate foundations, which renders any MBPP benchmark suboptimal for proper assessment.
@@
Table 3. Overview of competitive programming, code complexity, and code efficiency benchmarks.
@@
Competitive Programming CodeContests (Li et al., 2022a) 12 programming languages Avg. 203.7
APPS (Hendrycks et al., 2021a) Python Avg. 13.2
LiveCodeBench (Jain et al., 2024) Python Avg. 17.23
LeetCode (Tian et al., 2023) Python Avg. 135
CodeElo (Quan et al., 2025) N/A 408 problems
FVAPPS (Dougherty and Mehta, 2025) Python 4715 problems
KodCode (Xu et al., 2025b) 12 programming languages Avg. 7.52
FC2Code (Liu et al., 2022) Python 320 flowcharts
CodeContests+ (Wang et al., 2025b) Python 11,690 problems
LiveCodeBenchPro (Zheng et al., 2025a) Python 584 problems
Code Complexity CoRCoD (Sikka et al., 2019) Java 932
GeeksForGeeks (GFG) (Moudgalya et al., 2023) C++, Python ±1,400 per lang.&categ
@@
CodeComplex (Baik et al., 2024) Java, Python 4,900 per language
PythonSaga (Yadav and Singh, 2024) Python 185
Code Efficiency EffiBench (Huang et al., 2024) Python Self-defined, avg. 100
CODAL (Weyssow et al., 2024) Python 3 ref. / problem
PIE (Shypula et al., 2024) C++ 82.5(median, train)
COFFE (Peng et al., 2025) Python 756
Table 4. Overview of data science & domain-specific benchmarks. (JN refers to Jupyter Notebooks.)
@@
DS-1000 (Lai et al., 2022) Python Avg. 1.6 7 DS/ML libraries
NumpyEval (Zan et al., 2022b) Python Avg. 20 functions NumPy (101 problems)
@@
PandasEval (Zan et al., 2022b) Python Avg. 20 functions Pandas (101 problems)
@@
JuICe (Agashe et al., 2019) Python, JN N/A Cell completion
@@
DSP (Chandel et al., 2022) Python, JN Available Cell completion
@@
ExeDS (Huang et al., 2022) Python, JN Execution Based Cell generation
@@
DSEval (Zhang et al., 2024a) Python custom appraoch Models Evaluated via the DSEval
@@
TorchDataEval (Zan et al., 2022a) Python 50 Private PyTorch data library
MonkeyEval (Zan et al., 2022a) Python 101 Private Pandas library fork
BeatNumEval (Zan et al., 2022a) Python 101 Private NumPy variant
Bio-Coder (Tang et al., 2024b) Python 1,026 Identify and import necessary
@@
Bio-Coder-Rosalind (Tang et al., 2024b) Python 253 golden solution Generate code for question
WebApp1k (Cui, 2024) React Available evaluates whether a model can
@@
Table 5. Overview of Mathematical Reasoning Benchmarks.
@@
MATH (Hendrycks et al., 2021b) English 12,500
MATH500 (Lightman et al., 2023) English 500
MathQA (Amini et al., 2019) English 37,297
MathQA-Python (Austin et al., 2021) Python 23,914
MathQA-X (Athiwaratkun et al., 2022) Python, Java, JS 1,883 per language
LiLa(Mishra et al., 2022) Python 133,815 questions, 358,769 programs
MultiArith (Roy and Roth, 2015) English 600
GSM8K (Gao et al., 2022) English 1,320
GSM-HARD (Gao et al., 2022) English 1,320
TheoremQA (Chen et al., 2023) English 800
PECC (Haller et al., 2024) Python 1,006
BRIGHT (Su et al., 2024) English 395
@@
Table 6. Overview of Natural Language Benchmarks.
Category Name Language(s) No. of Problems
Text2Code CoNaLa (Yin et al., 2018) English $\rightarrow$ Python 2,879
MCoNaLa (Wang et al., 2022) {Spanish, Japanese, 896

### v3

@@
Abstract
Benchmarks are essential for unified evaluation and reproducibility. The rapid rise of Artificial Intelligence for Software Engineering (AI4SE) has produced numerous benchmarks for tasks such as code generation and bug repair. However, this proliferation has led to major challenges: (1) fragmented knowledge across tasks, (2) difficulty in selecting contextually relevant benchmarks, (3) lack of standardization in benchmark creation, and (4) flaws that limit utility. Addressing these requires a dual approach: systematically mapping existing benchmarks for informed selection and defining unified guidelines for robust, adaptable benchmark development. We conduct a review of 247 studies, identifying 273 AI4SE benchmarks since 2014. We categorize them, analyze limitations, and expose gaps in current practices. Building on these insights, we introduce BenchScout, an extensible semantic search tool for locating suitable benchmarks. BenchScout employs automated clustering with contextual embeddings of benchmark-related studies, followed by dimensionality reduction. In a user study with 22 participants, BenchScout achieved usability, effectiveness, and intuitiveness scores of 4.5, 4.0, and 4.1 out of 5. To improve benchmarking standards, we propose BenchFrame, a unified approach to improve benchmark quality. Applying BenchFrame to HumanEval yielded HumanEvalNext, which features corrected errors, improved language conversion, higher test coverage, and greater difficulty. Evaluating 10 state-of-the-art code models on HumanEval, HumanEvalPlus, and HumanEvalNext revealed average pass-at-1 drops of 31.22% and 19.94%, respectively, underscoring the need for continuous benchmark refinement. We further examine BenchFrame’s scalability through an agentic pipeline and confirm its generalizability on the MBPP dataset. Lastly, we publicly release the material of our review, user study, and the enhanced benchmark.  [^1] https://github.com/AISE-TUDelft/AI4SE-benchmarks
I Introduction
@@
Figure 1: ChatGPT repeats the same HumanEval error (captured on Dec 2023)
A notable example of benchmark limitations in code generation evaluation is HumanEval [1], a widely-used dataset for assessing large language models (LLMs) such as Codex [1], Gemini [2], and GPT-4 [3]. HumanEval was downloaded 82 thousand times in July 2025 on a single platform (Hugging Face)  [^2] https://huggingface.co/datasets/openai/openai_humaneval, reflecting the strong and sustained interest in this benchmark. Despite its broad adoption, HumanEval contains numerous flaws and inconsistencies [4]. For instance, Task 47, which requires computing the median of a numerical list, incorrectly states the median of the list [-10,4,6,1000,10,20] is $15$. When queried on this task, ChatGPT-3.5 Turbo reproduced the incorrect result (see Figure 1). This suggests potential data contamination and benchmark overfitting which can artificially inflate performance scores.
As HumanEval remains a widely-used benchmark in both AI and software engineering communities, several efforts have sought to expand its language support [5, 6, 7, 8] or improve test coverage [9, 10]. However, these extensions often build upon the original dataset without addressing its fundamental deficiencies, allowing inherent issues, such as flawed canonical solutions, vague problem definitions, incorrect tests, and insufficient coverage to persist. Moreover, LLM-augmented improvements, such as automatic translation, often lack rigorous quality control as well. Lastly, as models have advanced, HumanEval and similar popular benchmarks have become increasingly saturated, with close to 100% scores for recent models.  [^3] https://evalplus.github.io/leaderboard.html This necessitates continuous elevation of the complexity of programs to better reflect the capabilities of contemporary models.
@@
While BenchScout aids navigation, our analysis confirms that many foundational benchmarks are flawed, which means a systematic solution is required to deal with them. Based on the identified gaps and limitations in current benchmarks, we introduce BenchFrame, a peer-review-oriented methodology to improve the quality of the benchmarks of both existing and new benchmarks. To demonstrate its efficacy, we conduct a case study on HumanEval and present HumanEvalNext as an enhanced version.
When evaluating performance using HumanEval (original) and HumanEvalNext (our improved version based on the BenchFrame), we observe a substantial decline in pass@1 accuracy. Across ten state-of-the-art open-source code models, the average pass@1 score decreases by 31.2%, with a median drop of 26.0%. Performance remains significantly lower even on HumanEvalPlus [10], an enhanced version of HumanEval, with an average decline of $19.94\%$ in pass@1 scores. These results highlight the importance of continuously refining benchmarks to better guide future research and provide more realistic assessments of model performance.
@@
  • We conducted a comprehensive review of 247 studies, identified 273 AI4SE benchmarks, and analyzed their limitations and gaps (section III),
  • We developed and released BenchScout, an extensible semantic search tool to facilitate locating appropriate AI4SE benchmarks. Our user study with 22 participants demonstrated its effectiveness (section IV),
  • We propose a unified approach, BenchFrame, to improve the quality and reliability of benchmarks. A case study on HumanEval resulted in a refined, peer-reviewed benchmark, HumanEvalNext (section V). We assessed ten recent LLMs on this benchmark and verified significant drops in performance compared to the original HumanEval[1] and the enhanced version, HumanEvalPlus [10],
@@ II Research Questions
These research questions directly map to the three-stage approach of our study. RQ1 fulfills the diagnostic part of our objective by mapping the landscape and identifying its core problems. The findings from RQ1 motivate RQ2, which provides a practical tool to help the community navigate the benchmark fragmentation we uncovered. Finally, RQ3 addresses the root cause of the quality issues found in RQ1 by proposing and validating a systematic approach for improvement, thereby fulfilling the final part of our research objective.
III Existing Benchmarks, a Review
III-A Search Criteria and Quality Assessment
@@
We employed a systematic method to search for, identify, and classify AI4SE benchmarks. This procedure involved three primary stages: conducting structured searches, verifying credibility, and developing a taxonomy. We searched on two platforms, namely Google Scholar and Semantic Scholar, using the following keywords: “Benchmarks”, “Software Engineering”, “Large Language Models”, “Evaluation”, and “AI4SE”. We additionally searched for benchmarks present in the PapersWithCode datasets collection due to the popularity and wide usage of the platform.  [^6] This platform has since been discontinued and is no longer accessible. The previously used URL was https://paperswithcode.com/datasets
@@
We selected these keywords based on a preliminary assessment of highly-cited benchmark research. Three authors worked together to revise the search criteria to ensure they were both relevant and complete. Our selection criteria targeted primary studies published in English from 2014 to 2025. After the initial paper collection process, duplicates were removed. Two authors reviewed the relevance of the identified pieces of literature, during which the originality (that is, the status as primary study), reproducibility, and accessibility of each study were accessed.
@@
More specifically, for originality, we assessed whether a study was a primary study that introduced a new benchmark or a significant extension of an existing one. This filter was set in place to distinguish between benchmark-proposing papers from studies that merely applied them with a certain level preprocessing. In terms of Reproducibility and Accessibility, we evaluated whether the benchmark and its associated resources were made publicly available. We looked for the presence of a public repository (e.g., GitHub or HuggingFace) that would allow the community to access and use the benchmark. After this phase, we performed forward and backward snowballing. A total of 247 papers were found through this process.
@@
After the papers were collected, the authors worked together to develop and consistently improve a taxonomy to efficiently categorize the benchmarks and extract metadata. Through continuous discussions, we identified initial essential categories while systematically gathering additional details for each study, including DOI and publication date. An iterative strategy was used to design the taxonomy and categorize SE tasks, starting with overarching categories such as reasoning, synthesis, and debugging. When the categories became too broad or lacked cohesion, the authors refined them into more detailed subcategories, establishing a multi-tiered hierarchy. Furthermore, it is important to highlight that during the categorization process, the two authors involved discovered 10 instances of disagreement over category classifications, resulting in an agreement rate of 96.4%. This outcome can be attributed to the relatively objective nature of benchmark categorization. Many of the papers either clearly specify their intended task or imply it within the text, with the latter being the source of discrepancies in the few instances where the raters disagreed.
@@
In our evaluation, we not only considered benchmark datasets from academic sources, but also included those suggested by industry and subsequently adopted by scholars (for example, IBM CODAIT and Aider [11]). Our replication package also serves as a dynamic repository, allowing researchers to contribute additional benchmarks and related papers by submitting a pull request that includes the paper’s DOI.
@@
III-B Results of the Review
Figure 2: Number of published benchmark papers (Jan 2014-Aug 2025).
@@ III-B Results of the Review
Additionally, upon analyzing the distribution of the tasks included in the review, we find that a great portion of benchmarks are related to Code Generation, more specifically, 34.4% of all the benchmarks included. Along with Code Understanding and Repair & Maintenance (with respectively 17.7% and 15.4%), these make up more than 67% of all benchmarks included. This, in itself, underscores a lack of attention to tasks such as vulnerability detection and code retrieval/search (cf. Figure 3).
The data reveal a notable change over time in the distribution of research categories throughout the observed period (cf. Figure 4). A key insight is the shift from code repair and maintenance to automated synthesis, marked by two opposing trends. Initially dominant, the Repair & Maintenance category gradually experiences a notable decline in its share. In contrast, starting around 2020, the Code Generation category gains prominence, which coincides with the advancement of large-scale language models. In the recent years presented, there is an emerging trend of thematic diversification; as the focus on Code Generation recedes from its high, categories like Code Understanding and Retrieval & Search show a relative resurgence. This suggests a community-wide transition towards tackling additional challenges related to verification, integration, and the understanding of algorithmically generated outputs. In the following, we summarize key insights from our review.
Code Generation (34.4%) Code Understanding (17.7%) Repair & Maintenance (15.4%) Other (13.4%) Math & Data (10.7%) Vulnerability Detection (5.4%) Retrieval & Search (3.0%)
Figure 3: High-level category distribution of included benchmarks (updated, alternating orientation).
Figure 4: Task distribution progression for the benchmarks included in the study.
TABLE I: Overview of AI4SE benchmarks stemming from HumanEval [1] (PL denotes Programming Language, NL denotes Natural Language).
@@
Original HumanEval [1] Python Avg. 7.7
Improved Language Support MultiPL-HumanEval [6] 18 PLs Avg. 7.7
HumanEval-Fix [7] 6 PLs Avg. 7.7
HumanEval-Explain [7] 6 PLs Avg. 7.7
HumanEval-Synthesize [7] 6 PLs Avg. 7.7
HumanEval-X [8] 5 PLs Avg. 7.7
Multi-HumanEval [5] 12 PLs Avg. 7.7
HumanEvalXL [12] 12 PLs, 23 NLs Avg. 8.33
mHumanEval [13] Python, 204 NLs Avg. 7.7
Improved Testing HumanEval+ [10] Python Scaled ×80
HumanEval-MINI [10] Python Scaled ×47
HE-Eval [9] Python Scaled ×14
@@
Extended EvoEval [14] Python Multiple categories,
@@
HumanEval Benchmark Family: Currently, one of the most popular AI4SE benchmarks is HumanEval [1], used to evaluate the performance of many notable code-aware models (e.g., Codex [1], Gemini [2], and GPT-4 [3]). This benchmark is used mainly for code synthesis, though there exist some variations for code repair and code explanation [7]. Table I presents the family of HumanEval benchmarks. After an in-depth analysis of these benchmarks, we identified the following issues: (1) incorrect tests, (2) lack of proper test coverage, (3) incorrect canonical solutions, and (4) imprecise problem definitions. While there are versions that have improved the language support [5, 6, 7, 8] and test coverage [9, 10], there is no version that contains all the improvements combined nor fixed the original issues. The issues for enhancing the original dataset can be generalized as follows:
@@
TABLE II: Overview of AI4SE benchmarks derived from MBPP [15].
@@
Original MBPP [15] Python 974
Improved Language MultiPL-MBPP [6] 18 PLs 354-397 per PL
Support MBXP [5] 13 PLs 848-974 per PL
Improved MBPP+ [10] Python 427
Testing MBPP-Eval [9] Python 974
MBPP Benchmark Family: Another AI4SE benchmark, highly similar in style and popularity compared to HumanEval, is MBPP [15]: Mostly Basic Python Problems. It contains nearly a thousand crowdsourced problems, where almost half of it is sanitized and separately released. Furthermore, several enhancements have been published for MBPP (Table II). Upon a more in-depth analysis of MBPP and its family of benchmarks, there are many signs suggesting deficient quality. One notable problem is the lack of proper testing, as MBPP originally only has three (rather trivial) tests per problem, which are all revealed in the prompt as well. With such a test suite in place, evaluation metrics become unstable and insignificant for proper comparison. The strength of the written tests and solutions themselves is not only troublesome in the original data but also the sanitized data features many flaws (even in corrected variants [10]). From negligible observations such as poor syntax (e.g., too many spaces, Python method names starting with a capital – this is a common convention to only use for classes) to uncaught bugs and edge cases that break the implementation. While there are enhancements that improve the language support and extend the test cases, they are all built upon inadequate foundations, which renders any MBPP benchmark suboptimal for proper assessment.
@@
TABLE III: Overview of competitive programming, code complexity, and code efficiency benchmarks.
@@
Competitive Programming CodeContests [16] 12 PLs Avg. 203.7
APPS [17] Python Avg. 13.2
LiveCodeBench [18] Python Avg. 17.23
LeetCode [19] Python Avg. 135
CodeElo [20] N/A 408 problems
FVAPPS [21] Python 4715 problems
KodCode [22] 12 PLs Avg. 7.52
FC2Code [23] Python 320 flowcharts
CodeContests+ [24] Python 11,690 problems
LiveCodeBenchPro [25] Python 584 problems
Code Complexity CoRCoD [26] Java 932
GeeksForGeeks (GFG) [27] C++, Python ±1,400 per lang.&categ
@@
CodeComplex [28] Java, Python 4,900 per language
PythonSaga [29] Python 185
Code Efficiency EffiBench [30] Python Self-defined, avg. 100
CODAL [31] Python 3 ref. / problem
PIE [32] C++ 82.5(median, train)
COFFE [33] Python 756
TABLE IV: Overview of data science & domain-specific benchmarks. (JN refers to Jupyter Notebooks.)
@@
DS-1000 [34] Python Avg. 1.6 7 DS/ML libraries
NumpyEval [35] Python Avg. 20 functions NumPy (101 problems)
@@
PandasEval [35] Python Avg. 20 functions Pandas (101 problems)
@@
JuICe [36] Python, JN N/A Cell completion
@@
DSP [37] Python, JN Available Cell completion
@@
ExeDS [38] Python, JN Execution Based Cell generation
@@
DSEval [39] Python custom appraoch Models Evaluated via the DSEval
@@
TorchDataEval [40] Python 50 Private PyTorch data library
MonkeyEval [40] Python 101 Private Pandas library fork
BeatNumEval [40] Python 101 Private NumPy variant
Bio-Coder [41] Python 1,026 Identify and import necessary
@@
Bio-Coder-Rosalind [41] Python 253 golden solution Generate code for question
WebApp1k [42] React Available evaluates whether a model can
@@
TABLE V: Overview of Mathematical Reasoning Benchmarks.
@@
MATH [43] English 12,500
MATH500 [44] English 500
MathQA [45] English 37,297
MathQA-Python [15] Python 23,914
MathQA-X [5] Python, Java, JS 1,883 per language
LiLa[46] Python 133,815 questions
358,769 programs
MultiArith [47] English 600
GSM8K [48] English 1,320
GSM-HARD [48] English 1,320
TheoremQA [49] English 800
PECC [50] Python 1,006
BRIGHT [51] English 395
@@
TABLE VI: Overview of Natural Language Benchmarks. These are specifically the Text2Code Bencharmks.
Name Language(s) # Problems
CoNaLa [52] English $\rightarrow$ Python 2,879
MCoNaLa [53] {Spanish, Japanese, 896